A triangular lamina in the xy -plane such that its vertices are (0,0), (0,1) and (1,0). Suppose that the density function of the lamina is defined by p(x,y)=30xy gram per cubic centimetre. What is the total mass of the lamina and the center of gravity. The moment about the y-axis of the lamina.the moment about the x axis of the lamina


1
Expert's answer
2021-09-12T23:52:32-0400

Let "A=(0,1), B=(1, 0), O=(0,0)."

Line OA: "x=0, 0\leq y\leq1"

LineAB: "y=-x+1, 0\leq x\leq1"

Line OB: "y=0, 0\leq x\leq 1"


Given "\rho(x,y)=30xy"

Find the mass of the lamina


"m=\iint_D\rho(x,y)dA=\displaystyle\int_{0}^1\displaystyle\int_{0}^{1-x}30xydydx=""=30\displaystyle\int_{0}^1x\big[{y^2\over 2}\big]\begin{matrix}
 1-x \\
 0 
\end{matrix}dx=15\displaystyle\int_{0}^1(x-2x^2+x^3 )dx=""=15\big[{x^2 \over 2}-{2x^3 \over 3}+{x^4 \over 4}\big]\begin{matrix}
 1 \\
 0
\end{matrix}={5 \over 4} (units\ of\ mass)"

Mass of the lamina is "\dfrac{5}{4}" units of mass.


The moment of the lamina about the x-axis


"M_x=\iint_Dy\rho(x,y)dA=30\displaystyle\int_{0}^1\displaystyle\int_{0}^{1-x}xy^2dydx="




"=30\displaystyle\int_{0}^1x\big[{y^3\over 3}\big]\begin{matrix}
 1-x \\
 0 
\end{matrix}dx"




"=10\displaystyle\int_{0}^1(x-3x^2+3x^3-x^4 )dx"




"=10\big[{x^2 \over 2}-{3x^3 \over 3}+{3x^4 \over 4}-{x^5 \over 5}\big]\begin{matrix}
 1 \\
 0
\end{matrix}"

"=5-10+\dfrac{15}{2}-2={1 \over2}"

The moment of the lamina about the y-axis


"M_y=\iint_Dx\rho(x,y)dA=\displaystyle\int_{0}^1\displaystyle\int_{0}^{1-x}30x^2ydydx"




"=30\displaystyle\int_{0}^1x^2\big[{y^2\over 2}\big]\begin{matrix}
 1-x \\
 0 
\end{matrix}dx"




"=15\displaystyle\int_{0}^1(x^2-2x^3+x^4 )dx"




"=15\big[{x^3 \over 3}-{2x^4 \over 4}+{x^5 \over 5}\big]\begin{matrix}
 1 \\
 0
\end{matrix}"


"=5-\dfrac{15}{2}+3=\dfrac{1}{2}"

"M_x=\dfrac{1}{2}, M_y=\dfrac{1}{2}"

Find the coordinates of the center of mass


"\bar{x}={M_y\over m}=\dfrac{\dfrac{1}{2}}{\dfrac{5}{4}}=\dfrac{2}{5}"



"\bar{y}={M_x\over m}=\dfrac{\dfrac{1}{2}}{\dfrac{5}{4}}=\dfrac{2}{5}"



Center of gravity is"\bigg(\dfrac{2}{5},\dfrac{2}{5}\bigg)."